OpenStudy (anonymous):

Algebra 2---Please help with this problem!!! Will give medals and hugs! Show that the sum of the reciprocals of three different positive integers is greater than 6 times the reciprocal of their product.

OpenStudy (anonymous):

You have \[\frac{1}{a} + \frac{1}{b}+\frac{1}{c} > \frac{6}{abc}\] Add the fractions on the left first

OpenStudy (anonymous):

3/abc ?

OpenStudy (anonymous):

How did you get 3 in the numerator? Get common denominators on each of the fractions then add them

OpenStudy (anonymous):

How do you do that with letters?

OpenStudy (anonymous):

They are variables for the three different positive integers. You don't evaluate them.

OpenStudy (anonymous):

So you could just replace it with 1,2,3 ?

OpenStudy (anonymous):

You want to prove it for any three numbers so you leave them as variables.

OpenStudy (anonymous):

How do you do that?

OpenStudy (anonymous):

\[\frac{1}{a}+\frac{1}{b}+\frac{1}{c}=\frac{ab+ac+bc}{abc}\]

OpenStudy (anonymous):

do your real job is to show that \(ab +ac+bc>6\) which is easy enough since the smallest possible different positive integers are 1, 2, and 3

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

So what is the answer? I'm confused on how you get that.

OpenStudy (anonymous):

first of all i added up the fractions

OpenStudy (anonymous):

here lets do it this way what are the three smallest possible positive integers?

OpenStudy (anonymous):

1,2, and 3 right?

OpenStudy (anonymous):

@satellite73